# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 5.3 Edges
The next theorem says that, in minimal Taylor algebras, every “thick” edge, in the terminology of [26, 27], is automatically a subuniverse. This property is a simple consequence of the results we have already stated, whereas it was relatively painful to achieve using the original approach. We additionally obtain that semilattice and majority edges have unique witnessing congruences.
@@
(link to proof) Let $(a,b)$ be an edge (semilattice, majority, or abelian) of a minimal Taylor algebra $\mathbf{A}$ and $\theta$ a witnessing congruence of $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$.
@@
    If $(a,b)$ is a semilattice edge, then $\mathbf{E}/\theta$ is term equivalent to a two-element semilattice with absorbing element $b/\theta$.
@@
    If $(a,b)$ is a majority edge, then $\mathbf{E}/\theta$ is term equivalent to a two-element majority algebra.
@@
    if $(a,b)$ is an abelian edge, then $\mathbf{E}/\theta$ is term equivalent to an affine Mal’cev algebra of an abelian group isomorphic to $\mathbb{Z}/{m}$ for some positive integer $m$, where $\mathbb{Z}/{m}$ denotes the group of integers modulo $m$.
Moreover a semilatice edge is witnessed by exactly one congruence of $\mathbf{E}$, and that congruence is maximal. The same holds for majority edges.
For minimal edges we can say a bit more. If $(a,b)$ is a minimal edge witnessed by $\theta$, a congruence on $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$, then $\mathbf{E}/\theta$ is simple. In particular, for abelian minimal edges, $\mathbf{E}/\theta$ is an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. Moreover, such an $\mathbf{E}$ has a unique maximal congruence as shown in the next proposition. This implies that the type of a minimal edge is unique and so is the direction of a semilattice minimal edge and the prime $p$ associated to an abelian minimal edge.
@@
(link to proof) Let $(a,b)$ be a minimal edge in a minimal Taylor algebra. Then $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$ has a unique maximal congruence equal to $\mu_{\mathbf{E}}$. In particular, minimal edges have unique types.
@@
(link to proof) Let $(a,b)$ be a minimal semilattice edge in a minimal Taylor algebra. Then $\{a,b\}$ is a subuniverse of $\mathbf{A}$, so $\operatorname{Sg}_{\mathbf{A}}(a,b)=\{a,b\}$ and the witnessing congruence is the equality.
Unfortunately, majority and abelian edges do not simplify in a similar way; see Example 5.16 and Example 5.17. Weaker versions of Proposition 5.15 have been developed by Bulatov (comp. Lemma 12 and Corollary 13 in [26]) to deal with this problem.
@@ 5.4 Absorption and edges
Let $\mathbf{A}$ be an algebra, let $B\subseteq A$ and let $(b,a)$ be an edge. We say that $B$ is stable under $(b,a)$ if, for every witnessing congruence $\theta$ of $\operatorname{Sg}_{\mathbf{A}}(b,a)$ such that $b/\theta$ intersects $B$, each $\theta$-class intersects $B$.
@@
(link to verification) We consider the four-element algebra $\mathbf{A}=(\{0,1,2,*\},\cdot)$ with binary operation normal-$\cdot$ given by
$\cdot$ 00 $1$ $2$**
00 00 $2$ $1$**
@@
For absorption of higher arity the connection to edges is not as tight as for 2-absorption. Nevertheless, one direction still works and both directions work for singletons.
@@
(link to proof) Any absorbing set of a minimal Taylor algebra $\mathbf{A}$ is stable under semilattice and abelian edges. Moreover, for any $b\in A$ the following are equivalent.
@@
(link to proof) Every minimal Taylor algebra $\mathbf{A}$ has a ternary term operation $f$ such that if $(a,b)$ is an edge witnessed by $\theta$ on $\mathbf{E}=\operatorname{Sg}_{\mathbf{A}}(a,b)$, then
  • if $(a,b)$ is a semilattice edge, then $f(x,y,z)=x\vee y\vee z$ on $\mathbf{E}/\theta$ (where $b/\theta$ is the top);
  • if $(a,b)$ is a majority edge, then $f$ is the majority operation on $\mathbf{E}/\theta$ (which has two elements);
  • if $(a,b)$ is an abelian edge, then $f(x,y,z)=x-y+z$ on $\mathbf{E}/\theta$;
  • $f$ witnesses all the 3-absorptions $B\trianglelefteq_{3}\mathbf{A}$;
  • any binary operation obtained from $f$ by identifying two arguments witnesses all the 2-absorptions $B\trianglelefteq_{2}\mathbf{A}$.
@@
(link to proof) If $\mathbf{A}$ is a minimal Taylor algebra, then $\operatorname{Clo}(A;f)=\operatorname{Clo}(\mathbf{A})$ for any operation $f$ satisfying the first three items in Theorem 5.23.
@@ 6 Omitting types
It turns out that within minimal Taylor algebras these “omitting types” conditions are often equivalent to important properties of algebras. In the theorems below we prove the equivalence of the following four types of conditions: (i) the absence of edges of a certain type (equivalently, minimal edges of the same type); (ii) properties of absorption and the four types in Zhuk’s approach; (iii) the existence of a certain special term operations; (iv) algorithmic properties of the CSP. Theorems in this section are consequences of the theory we have already built in the previous section and known results (see [9]). The proofs are in Section 10.
@@
    $\mathbf{A}$ has a wnu term operation of every arity $n\geq 3$.
@@
Minimal Taylor algebras omitting other types of edges do have significantly stronger properties than general Taylor algebras omitting those edges. Minimal $\mathrm{s}$-free algebras are exactly those for which option (a) in Theorem 3.7 does not hold, and that have few subpowers [27]. Few subpowers algebras can be characterized by the existence of an edge term operation [12] in general. In minimal Taylor algebras, the second strongest edge operation always exists – the 3-edge operation defined by the identities $e(y,y,x,x)=e(y,x,y,x)=e(x,x,x,y)=x$. This is significant, because the exponent in the running time of the few subpowers algorithm depends on the least $k$ such that the algebra has a $k$-edge term operation. The number $3$ here is best possible: a $2$-edge operation is the same as a Mal’cev operation appearing in Theorem 6.6.
@@
    Every center (3-absorbing subuniverse) of any $\mathbf{B}\leq\mathbf{A}$ 2-absorbs $\mathbf{B}$, i.e., case (b) implies case (a) in Theorem 3.7 in all the subalgebras of $\mathbf{A}$.
@@
Surprisingly, if along with $\mathrm{m}$-freeness we also limit the type of abelian edges allowed in an algebra, the resulting condition is equivalent to the existence of a binary commutative term operation. This is interesting to us (the authors), since we did not regard the existence of a commutative term operation to be a natural requirement for the CSP (cf. [9]). We call an abelian edge $(a,b)$ a $\mathbb{Z}/{2}$-edge if the corresponding affine Mal’cev algebra $\operatorname{Sg}(a,b)/\theta$ is isomorphic to the affine Mal’cev algebra of $\mathbb{Z}/{2}$.
@@
    $\operatorname{Clo}(\mathbf{A})$ can be generated by a collection of binary operations.
@@
Minimal Taylor $\mathrm{am}$-free algebras are exactly those which have wnu operations of every arity $n\geq 2$. These are exactly the minimal spirals in the terminology of [18] and a significant property is that for every $(a,b)$ such that neither $(a,b)$ nor $(b,a)$ is a minimal semilattice edge, there is a surjective homomorphism from $\operatorname{Sg}\{a,b\}$ onto the (three-element) free semilattice on two generators.
The $\mathrm{sm}$-free minimal Taylor algebras are those where cases (a) and (b) in Theorem 3.7 do not occur. Additionally, these are exactly the hereditarily absorption free algebras studied in [8] and, also, the algebras with a Mal’cev term operation – a type of operation that played a significant role in the CSP [21].
@@
Finally, the $\mathrm{as}$-free algebras are those where cases (a) and (c) in Theorem 3.7 do not occur and those that have bounded width and few subpowers. It is known [42, 12] that the latter property in general implies having a near-unanimity term operation of some arity. Surprisingly, in minimal Taylor algebras, the arity goes down directly to three. In the algorithmic language, these algebras have strict width two [38, 9].
@@
    $\mathbf{A}$ has a near unanimity term operation.
@@
    $\mathbf{A}$ has a majority term operation.
PART II: Technical details
@@ 7.2 Algebras
Algebras, i.e. structures with a purely functional signature, will be denoted by boldface capital letters (e.g., $\mathbf{A}$) and their universes (also called domains) typically by the same letter in the plain font (e.g., $A$). The basic general algebraic concepts, such as subuniverses, subalgebras, products, and quotients modulo congruences are used in the standard way (see, e.g. [11]). An algebra is nontrivial if it has at least two elements, otherwise it is trivial. We use $B\leq\mathbf{A}$ to mean that $B$ is a subuniverse of $\mathbf{A}$. We often abuse notation and write $t$ both for a term and the induced term operation $t^{\mathbf{A}}$ of an algebra $\mathbf{A}$. By a subpower we mean a subuniverse (or a subalgebra) of a finite power. Subpowers are the same as invariant relations and we may also call them compatible relations. The set of all subpowers is denoted $\operatorname{Inv}(\mathbf{A})$. The subuniverse (or the subalgebra) of $\mathbf{A}$ generated by a set $X\subseteq A$ is denoted $\operatorname{Sg}_{\mathbf{A}}(X)$ or $\operatorname{Sg}_{\mathbf{A}}(x_{1},\dots,x_{n})$ when $X=\{x_{1},\dots,x_{n}\}$. An algebra is simple if it has only the trivial congruences – the equality relation and the full relation.
@@
An operation $f$ on $A$ is a semilattice operation if it is binary, commutative, idempotent, and associative; cyclic operation if $f(a_{1},\dots,a_{n})=f(a_{2},\dots,a_{n},a_{1})$ for all ${\mathbf{a}}\in A^{n}$; weak near unanimity (wnu) operation if $f(b,a,\dots,a)=f(a,b,a,\dots,a)=\dots=f(a,\dots,a,b)$ for all $a,b\in A$; near unanimity operation if its a wnu and $f(b,a,\dots,a)=a$; a majority operation if it is a ternany near unanimity operation; a Mal’cev operation if $f(b,a,a)=f(a,a,b)=b$ for all $a,b\in A$; and affine Mal’cev operation if $f(a,b,c)=a-b+c$ for all $a,b,c\in A$, where +,+,-+ , - are computed with respect to an abelian group structure on $A$. For an odd $n$, the $n$-ary majority operation on a two-element domain is denoted $\operatorname{maj}_{n}$, i.e., $\operatorname{maj}_{n}(a_{1},\dots,a_{n})=a$ iff the majority of the $a_{i}$ is $a$. A semilattice is an algebra with a semilattice operation (and no other operations); an affine Mal’cev algebra is an algebra with an affine Mal’cev operation; and an affine algebra is an algebra whose term operations are exactly the idempotent term operations of a module over a unital ring.
A (function) clone is a set of operations $\mathscr{C}$ on a set $A$ which contains all the projections $\operatorname{proj}^{n}_{i}$ (the $n$-ary projection to the $i$-th coordinate, i.e., $\operatorname{proj}^{n}_{i}({\mathbf{a}})=a_{i}$) and is closed under composition, i.e., $f(g_{1},\dots,g_{n})\in\mathscr{C}$ whenever $f\in\mathscr{C}$ is $n$-ary and $g_{1}$, …, $g_{n}\in\mathscr{C}$ are all $m$-ary, where $f(g_{1},\dots,g_{n})$ denotes the operation defined by $f(g_{1}(x_{1},\dots,x_{m}),\dots,g_{n}(x_{1},\dots,x_{m}))$. By $\operatorname{Clo}(\mathbf{A})$ ($\operatorname{Clo}_{n}(\mathbf{A})$, respectively), we denote the clone of all term operations (the set of all $n$-ary term operations, respectively) of $\mathbf{A}$. An algebra $\mathbf{B}$ is a reduct of $\mathbf{A}$ if they have the same universe $A=B$ and $\operatorname{Clo}(\mathbf{B})\subseteq\operatorname{Clo}(\mathbf{A})$. Algebras $\mathbf{A}$ and $\mathbf{B}$ are term-equivalent if each of them is a reduct of the other, i.e., $\operatorname{Clo}(\mathbf{A})=\operatorname{Clo}(\mathbf{B})$.
An algebra $\mathbf{A}$ is polynomially complete if every operation on $A$ is in the clone generated by $\mathbf{A}$ together with the constant operations. Note that $\mathbf{A}$ is polynomially complete if and only if it has no proper reflexive irredundant subpowers.
@@ 7.4 Subpowers and pp-definitions
We have already mentioned in Subsection 7.1 that the relational composition $R+S$, the inverse relation $-R$, and left and right centers of a binary relation $R$ are pp-definable from $\{R,S\}$. Therefore if $R$ and $S$ are subuniverses of $\mathbf{A}^{2}$, then $R+S$ is a subuniverse of $\mathbf{A}^{2}$, and $-R$ as well as the centers of $R$ are subuniverses of $\mathbf{A}$ (and this also applies to the multisorted setting when $R$ and $S$ are subuniverses of $\mathbf{A}\times\mathbf{B}$). We now discuss further such observations that we use often in this paper.
If $R$ is a subuniverse of $\mathbf{A}^{n}$, then the projection onto a set of coordinates $I\subseteq\{1,2,\dots,n\}$ is a subpower of $\mathbf{A}$ as well. We sometimes fix some coordinate $i$ to a subuniverse $B\leq\mathbf{A}$ before projecting, i.e., we consider the relation $R^{\prime}(x_{1},\dots,x_{n})\equiv R(x_{1},\dots,x_{n})\wedge B(x_{i})$.
Let $R$ now be a subdirect subuniverse of $\mathbf{A}^{2}$. Recall that $R$ is linked if $S=(R-R)+\dots+(R-R)$ is equal to $A^{2}$ when we take a sufficiently large number of summands. In general, $S$ is an equivalence relation on $A$ which is a subuniverse of $\mathbf{A}^{2}$ – a congruence of $\mathbf{A}$. In particular, if $\mathbf{A}$ is simple, then $S$ is either $A^{2}$ (so $R$ is linked) or the equality relation, in which case $R$ is a graph of a permutation $A\to A$ – an automorphism $\mathbf{A}\to\mathbf{A}$. So, for a simple algebra $\mathbf{A}$, an irredundant binary relation $R\leq_{sd}\mathbf{A}^{2}$ is necessarily linked.
One type of subpower is of particular importance. Fix $n$ and consider the subalgebra $\mathbf{F}$ of $\mathbf{A}^{A^{n}}$ with universe $F=\operatorname{Clo}_{n}(\mathbf{A})$, the set of $n$-ary operations on $A$. Thus tuples in $F$ are $n$-ary term operations of $\mathbf{A}$, a coordinate of a tuple in $F$ is an element ${\mathbf{a}}$ of $A^{n}$, and the ${\mathbf{a}}$-th component of a tuple $f\in F$ is the value $f({\mathbf{a}})$. The algebra $\mathbf{F}$ is isomorphic to the free algebra for $\mathbf{A}$ over an $n$-element set of generators.

### v6

@@ 5.3 Edges
The next theorem says that, in minimal Taylor algebras, every “thick” edge, in the terminology of [Bulatov16:graph, Bulatov16:restricted], is automatically a subuniverse. This property is a simple consequence of the results we have already stated, whereas it was relatively painful to achieve using the original approach. We additionally obtain that semilattice and majority edges have unique witnessing congruences.
@@
(link to proof) Let $(a,b)$ be an edge (semilattice, majority, or abelian) of a minimal Taylor algebra and $\theta$ a witnessing congruence of .
@@
    If $(a,b)$ is a semilattice edge, then is term equivalent to a two-element semilattice with absorbing element $b/\theta$.
@@
    If $(a,b)$ is a majority edge, then is term equivalent to a two-element majority algebra.
@@
    if $(a,b)$ is an abelian edge, then is term equivalent to an affine Mal’cev algebra of an abelian group isomorphic to $\mathbb{Z}/{m}$ for some positive integer $m$, where $\mathbb{Z}/{m}$ denotes the group of integers modulo $m$.
Moreover, a semilatice edge is witnessed by exactly one congruence of , and that congruence is maximal. The same holds for majority edges.
For minimal edges we can say a bit more. If $(a,b)$ is a minimal edge witnessed by $\theta$, a congruence on , then is simple. In particular, for abelian minimal edges, is an affine Mal’cev algebra of a group isomorphic to $\mathbb{Z}/{p}$. Moreover, such an has a unique maximal congruence as shown in the next proposition. This implies that the type of a minimal edge is unique and so is the direction of a semilattice minimal edge and the prime $p$ associated to an abelian minimal edge.
@@
(link to proof) Let $(a,b)$ be a minimal edge in a minimal Taylor algebra. Then has a unique maximal congruence equal to $\mu$. In particular, minimal edges have unique types.
@@
(link to proof) Let $(a,b)$ be a minimal semilattice edge in a minimal Taylor algebra. Then $\{a,b\}$ is a subuniverse of , so $\operatorname{Sg}(a,b)=\{a,b\}$ and the witnessing congruence is the equality.
Unfortunately, majority and abelian edges do not simplify in a similar way; see Example 5.16 and Example 5.17. Weaker versions of Proposition 5.15 have been developed by Bulatov (comp. Lemma 12 and Corollary 13 in [Bulatov16:graph]) to deal with this problem.
@@ 5.4 Absorption and edges
Let be an algebra, let $B\subseteq A$ and let $(b,a)$ be an edge. We say that $B$ is stable under $(b,a)$ if, for every witnessing congruence $\theta$ of $\operatorname{Sg}(b,a)$ such that $b/\theta$ intersects $B$, each $\theta$-class intersects $B$.
@@
(link to verification) We consider the four-element algebra $\mathbf{A}=(\{0,1,2,*\},\cdot)$ with binary operation $\cdot$ given by
$\cdot$ 00 $1$ $2$
00 00 $2$ $1$
@@
For absorption of higher arity the connection to edges is not as tight as for 2-absorption. Nevertheless, one direction still works, and both directions work for singletons.
@@
(link to proof) Any absorbing set of a minimal Taylor algebra is stable under semilattice and abelian edges. Moreover, for any $b\in A$ the following are equivalent.
@@
(link to proof) Every minimal Taylor algebra has a ternary term operation $f$ such that if $(a,b)$ is an edge witnessed by $\theta$ on , then
  • if $(a,b)$ is a semilattice edge, then $f(x,y,z)=x\vee y\vee z$ on  (where $b/\theta$ is the top);
  • if $(a,b)$ is a majority edge, then $f$ is the majority operation on  (which has two elements);
  • if $(a,b)$ is an abelian edge, then $f(x,y,z)=x-y+z$ on ;
  • $f$ witnesses all the 3-absorptions $B\trianglelefteq_{3}$;
  • any binary operation obtained from $f$ by identifying two arguments witnesses all the 2-absorptions $B\trianglelefteq_{2}$.
@@
(link to proof) If $\mathbf{A}$ is a minimal Taylor algebra, then $\operatorname{Clo}(A;f)=\operatorname{Clo}($ for any operation $f$ satisfying the first three items in Theorem 5.23.
@@ 6 Omitting types
It turns out that within minimal Taylor algebras these “omitting types” conditions are often equivalent to important properties of algebras. In the theorems below we prove the equivalence of the following four types of conditions: (i) the absence of edges of a certain type (equivalently, minimal edges of the same type); (ii) properties of absorption and the four types in Zhuk’s approach; (iii) the existence of a certain special term operations; (iv) algorithmic properties of the CSP. Theorems in this section are consequences of the theory we have already built in the previous section and known results (see [Barto17:polymorphisms]). The proofs are in Section 10.
@@
    has a wnu term operation of every arity $n\geq 3$.
@@
Minimal Taylor algebras omitting other types of edges do have significantly stronger properties than general Taylor algebras omitting those edges. Minimal $\mathrm{s}$-free algebras are exactly those for which option (a) in Theorem 3.7 does not hold, and that have few subpowers [Bulatov16:restricted]. Few subpowers algebras can be characterized by the existence of an edge term operation [Berman10:varieties] in general. In minimal Taylor algebras, the second strongest edge operation always exists – the 3-edge operation defined by the identities $e(y,y,x,x)=e(y,x,y,x)=e(x,x,x,y)=x$. This is significant, because the exponent in the running time of the few subpowers algorithm depends on the least $k$ such that the algebra has a $k$-edge term operation. The number $3$ here is best possible: a $2$-edge operation is the same as a Mal’cev operation appearing in Theorem 6.6.
@@
    Every center (3-absorbing subuniverse) of any 2-absorbs , i.e., case (b) implies case (a) in Theorem 3.7 in all the subalgebras of .
@@
Surprisingly, if along with $\mathrm{m}$-freeness we also limit the type of abelian edges allowed in an algebra, the resulting condition is equivalent to the existence of a binary commutative term operation. This is interesting to us (the authors), since we did not regard the existence of a commutative term operation to be a natural requirement for the CSP (cf. [Barto17:polymorphisms]). We call an abelian edge $(a,b)$ a $\mathbb{Z}/{2}$-edge if the corresponding affine Mal’cev algebra $\operatorname{Sg}(a,b)/\theta$ is isomorphic to the affine Mal’cev algebra of $\mathbb{Z}/{2}$.
@@
    $\operatorname{Clo}($ can be generated by a collection of binary operations.
@@
Minimal Taylor $\mathrm{am}$-free algebras are exactly those which have wnu operations of every arity $n\geq 2$. These are exactly the minimal spirals in the terminology of [Brady20:bw] and a significant property is that for every $(a,b)$ such that neither $(a,b)$ nor $(b,a)$ is a minimal semilattice edge, there is a surjective homomorphism from $\operatorname{Sg}\{a,b\}$ onto the (three-element) free semilattice on two generators.
The $\mathrm{sm}$-free minimal Taylor algebras are those where cases (a) and (b) in Theorem 3.7 do not occur. Additionally, these are exactly the hereditarily absorption free algebras studied in [Barto15:malcev] and, also, the algebras with a Mal’cev term operation – a type of operation that played a significant role in the CSP [Bulatov06:maltsev].
@@
Finally, the $\mathrm{as}$-free algebras are those where cases (a) and (c) in Theorem 3.7 do not occur and those that have bounded width and few subpowers. It is known [hobby88:tct, Berman10:varieties] that the latter property in general implies having a near-unanimity term operation of some arity. Surprisingly, in minimal Taylor algebras, the arity goes down directly to three. In the algorithmic language, these algebras have strict width two [Feder98:monotone, Barto17:polymorphisms].
@@
    has a near unanimity term operation.
@@
    has a majority term operation.
Part II Technical details
@@ 7.2 Algebras
Algebras, i.e. structures with a purely functional signature, will be denoted by boldface capital letters (e.g., ) and their universes (also called domains) typically by the same letter in the plain font (e.g., $A$). The basic general algebraic concepts, such as subuniverses, subalgebras, products, and quotients modulo congruences are used in the standard way (see, e.g. [Bergman12:UAbook]). An algebra is nontrivial if it has at least two elements, otherwise it is trivial. We use $B\leq$ to mean that $B$ is a subuniverse of . We often abuse notation and write $t$ both for a term and the induced term operation $t$ of an algebra . By a subpower we mean a subuniverse (or a subalgebra) of a finite power. Subpowers are the same as invariant relations and we may also call them compatible relations. The set of all subpowers is denoted $\operatorname{Inv}($. The subuniverse (or the subalgebra) of $\mathbf{A}$ generated by a set $X\subseteq A$ is denoted $\operatorname{Sg}_{\mathbf{A}}(X)$ or $\operatorname{Sg}_{\mathbf{A}}(x_{1},\dots,x_{n})$ when $X=\{x_{1},\dots,x_{n}\}$. An algebra is simple if it has only the trivial congruences – the equality relation and the full relation.
@@
An operation $f$ on $A$ is a semilattice operation if it is binary, commutative, idempotent, and associative; cyclic operation if $f(a_{1},\dots,a_{n})=f(a_{2},\dots,a_{n},a_{1})$ for all ${\mathbf{a}}\in A^{n}$; weak near unanimity (wnu) operation if $f(b,a,\dots,a)=f(a,b,a,\dots,a)=\dots=f(a,\dots,a,b)$ for all $a,b\in A$; near unanimity operation if it is a wnu and $f(b,a,\dots,a)=a$; a majority operation if it is a ternary near unanimity operation; a Mal’cev operation if $f(b,a,a)=f(a,a,b)=b$ for all $a,b\in A$; and affine Mal’cev operation if $f(a,b,c)=a-b+c$ for all $a,b,c\in A$, where +,+,-+ , - are computed with respect to an abelian group structure on $A$. For an odd $n$, the $n$-ary majority operation on a two-element domain is denoted $\operatorname{maj}_{n}$, i.e., $\operatorname{maj}_{n}(a_{1},\dots,a_{n})=a$ iff the majority of the $a_{i}$ is $a$. A semilattice is an algebra with a semilattice operation (and no other operations); an affine Mal’cev algebra is an algebra with an affine Mal’cev operation; and an affine algebra is an algebra whose term operations are exactly the idempotent term operations of a module over a unital ring.
A (function) clone is a set of operations $\mathscr{C}$ on a set $A$ which contains all the projections $\operatorname{proj}^{n}_{i}$ (the $n$-ary projection to the $i$-th coordinate, i.e., $\operatorname{proj}^{n}_{i}({\mathbf{a}})=a_{i}$) and is closed under composition, i.e., $f(g_{1},\dots,g_{n})\in\mathscr{C}$ whenever $f\in\mathscr{C}$ is $n$-ary and $g_{1}$, …, $g_{n}\in\mathscr{C}$ are all $m$-ary, where $f(g_{1},\dots,g_{n})$ denotes the operation defined by $f(g_{1}(x_{1},\dots,x_{m}),\dots,g_{n}(x_{1},\dots,x_{m}))$. By $\operatorname{Clo}(\mathbf{A})$ ($\operatorname{Clo}_{n}(\mathbf{A})$, respectively), we denote the clone of all term operations (the set of all $n$-ary term operations, respectively) of $\mathbf{A}$. An algebra is a reduct of if they have the same universe $A=B$ and $\operatorname{Clo}($. Algebras and are term-equivalent if each of them is a reduct of the other, i.e., $\operatorname{Clo}($.
An algebra is polynomially complete if every operation on $A$ is in the clone generated by  together with the constant operations. Note that is polynomially complete if and only if it has no proper reflexive irredundant subpowers.
@@ 7.4 Subpowers and pp-definitions
We have already mentioned in Subsection 7.1 that the relational composition $R+S$, the inverse relation $-R$, and left and right centers of a binary relation $R$ are pp-definable from $\{R,S\}$. Therefore if $R$ and $S$ are subuniverses of , then $R+S$ is a subuniverse of , and $-R$ as well as the centers of $R$ are subuniverses of (and this also applies to the multisorted setting when $R$ and $S$ are subuniverses of ). We now discuss further such observations that we use often in this paper.
If $R$ is a subuniverse of , then the projection onto a set of coordinates $I\subseteq\{1,2,\dots,n\}$ is a subpower of as well. We sometimes fix some coordinate $i$ to a subuniverse $B\leq$ before projecting, i.e., we consider the relation $R^{\prime}(x_{1},\dots,x_{n})\equiv R(x_{1},\dots,x_{n})\wedge B(x_{i})$.
Let $R$ now be a subdirect subuniverse of . Recall that $R$ is linked if $S=(R-R)+\dots+(R-R)$ is equal to $A^{2}$ when we take a sufficiently large number of summands. In general, $S$ is an equivalence relation on $A$ which is a subuniverse of – a congruence of . In particular, if is simple, then $S$ is either $A^{2}$ (so $R$ is linked) or the equality relation, in which case $R$ is a graph of a permutation $A\to A$ – an automorphism . So, for a simple algebra , an irredundant binary relation $R\leq_{sd}$ is necessarily linked.
One type of subpower is of particular importance. Fix $n$ and consider the subalgebra of with universe $F=\operatorname{Clo}_{n}($, the set of $n$-ary operations on $A$. Thus tuples in $F$ are $n$-ary term operations of , a coordinate of a tuple in $F$ is an element ${\mathbf{a}}$ of $A^{n}$, and the ${\mathbf{a}}$-th component of a tuple $f\in F$ is the value $f({\mathbf{a}})$. The algebra is isomorphic to the free algebra for over an $n$-element set of generators.